OpenStudy (anonymous):

Write the inverse variation equation for the relationship: y varies inversely with x and y=8 and x=-2

OpenStudy (gorv):

y=c/x c=proportionality constant given y=8 x=-2 putting it in equation 8=c/(-2) 8*(-2)=c c=-16 y= -16/x

OpenStudy (gorv):

@smileykylee

OpenStudy (anonymous):

thanks

OpenStudy (gorv):

welcome @smileykylee